OpenStudy (blayne--moore):

What is the equation of a line passing through (-6, 5) and having a slope of ?

OpenStudy (blayne--moore):

the slope was 1 over 3

OpenStudy (blayne--moore):

idk

OpenStudy (kevin):

What's the slope?

OpenStudy (kevin):

1/3?

OpenStudy (saadsword):

y=mx+b m is the slope b is the y intercept y=1/3 x + b Now since you know the slope graph the equation and see wher it intersects the y axis

OpenStudy (saadsword):

and plug it in as b

OpenStudy (mathmate):

Or use the point-slope form of the equation: y-y1=m(x-x1) m=1/3, point P(x1,y1)=(-6,5)

OpenStudy (mathmale):

Since you are given the slope of this line and one point on the line, use the"point-slope form of the equation of a straight line:"\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

See mathmate's presentation above; it is correct. Finish this work, please. Simplify the resulting equation as much as possible. Prefered final form is y=mx+b.